# Commutative algebra inspired by modularity lifting

### v1

@@ Abstract
There are many excellent sources that give an overview of the work of Wiles’ proof; see, for example, [13, Chapter 1] and [14]. Roughly speaking, using Mazur’s [28] deformation theory of Galois representations, Wiles reduces the Shimura, Taniyama, Weil conjecture to proving that certain surjective maps $\varphi_{\Sigma}\colon R_{\Sigma}\to\mathbb{T}_{\Sigma}$ of (noetherian, commutative) local rings are isomorphisms. The ring $R_{\Sigma}$ parametrizes Galois deformations subject to constraints determined by a finite collection of primes $\Sigma$ and $\mathbb{T}_{\Sigma}$ is a Hecke algebra subject to the same constraints, which acts on a certain space of modular forms, $M_{\Sigma}$. To prove that $\varphi_{\Sigma}$ is an isomorphism, Wiles first proves that $M_{\Sigma}$ is free as a $\mathbb{T}_{\Sigma}$-module—this is a delicate argument—and also that $\mathbb{T}_{\Sigma}$ is a complete intersection. He then uses these properties to prove that $\varphi_{\Sigma}$ is an isomorphism; as corollary, he deduces the freeness of $M_{\Sigma}$ also as an $R_{\Sigma}$-module. It is Diamond’s insight [15] that one can prove the latter statement directly and that this implies $\varphi_{\Sigma}$ is an isomorphism of complete intersections, and also that $M_{\Sigma}$ is free over $\mathbb{T}_{\Sigma}$.
@@
where the map on the left is a homomorphism of rings, say with kernel $I$, and $M_{\varnothing}\cong M_{\infty}/IM_{\infty}$. The point of the patching process is that $R_{\infty}$ is a regular local ring and $M_{\infty}$ is a maximal Cohen-Macaulay module over it. Thus the Auslander-Buchsbaum formula implies that $M_{\infty}$ is free $R_{\infty}$-module, and hence that $M_{\varnothing}$ is a free $R_{\varnothing}$-module, as desired; see Proposition 2.1. It turns out that $I$ is generated by a regular sequence, so as a byproduct of the patching argument one gets that $R_{\varnothing}$ is a complete intersection; this is special to Wiles’ context and important for the passage from the minimal to the general, non-minimal, case. This passage is facilitated by a new criterion, discovered by Diamond [15] (building on work of Wiles and Lenstra [26]) that detects when a finitely generated module over a local ring is free and the ring is a complete intersection; see Theorem 3.1.
@@
Commutative algebra thus enters in both steps of the proof. As such, the result that Wiles and Diamond wanted (the Auslander-Buchsbaum formula) had been available since the late 1960’s. In contrast, to tackle the non-minimal case, Wiles and Diamond needed to develop new commutative algebra, and particular a numerical criterion (in terms that are meaningful in number theory) for detecting freeness of modules over local rings. Nevertheless, each step raises new questions, and lead to new results, in commutative algebra. This is what is described below.
@@
The patching method is (or so it seems to me) specific to the number theoretic context. In Section 2, I discuss the commutative algebra that is partly motivated by a desire to simplify, and even entirely do away with, the patching argument of Taylor and Wiles to establish modularity in the minimal case. It starts with a conjecture of de Smit (now a theorem of Brochard’s) that, in hindsight, can be interpreted as a converse to a well-known (to commutative algebraists!) result of Nagata’s on a special kind of surjective complete intersection maps [^1] However, the hypothesis of the de Smit conjecture provides the seed of a patching datum that develops into full fledged patching. The patching argument has become an indispensable tool in the number theorist’s toolkit: it introduces a useful rigidity and robustness, allowing one to, for example, invert the prime $p$ when proving modularity of $p$-adic Galois representations) which the hypothesis of de Smit’s conjecture does not provide..
@@
The patching method of Taylor and Wiles has been, and continues to be, developed extensively, and is now available also in the non-minimal case, thanks to work of Kisin [24], and this has often allowed one to bypass the numerical criterion. However, although this has lead to much more general modularity lifting theorems, among other application, it has been at the expense of some refined torsion information; all this is explained in the introduction to [19]. See also the discussion after Theorem 3.3. The main contribution of op. cit. is to establish a numerical criterion that applies in greater generality, so that it can be applied to rings and modules after patching.
@@
In what follows, I assume familiarity with basic concepts from commutative algebra; much of what is needed can be found in Chapter 1 of the book by Bruns and Herzog [8], and I take that as the standard reference. The homological algebra notions that come up are already present, for most parts, in the book by Cartan and Eilenberg [12]. Given a local ring $A$, I write ${\mathfrak{m}}_{A}$ for its maximal ideal and $k_{A}$ for its residue field $A/{\mathfrak{m}}_{A}$. The embedding dimension of A, denoted $\operatorname{edim}A$, is $k_{A}$ rank of the Zariski tangent space ${\mathfrak{m}}_{A}/{\mathfrak{m}}_{A}^{2}$; equivalently, the minimal number of generators of the ideal ${\mathfrak{m}}_{A}$.
@@
Subsequent to Diamond’s work, Bart de Smit conjectured that if $A\to B$ is a local homomorphism of artinian local rings with $\operatorname{edim}A=\operatorname{edim}B$, then any $B$-module that is flat as an $A$-module is also flat as a $B$-module. This strengthens Proposition 2.1 in the case $A$ and $B$ have the same embedding dimension, and allows one to dispense with patching technique in the work of Wiles; see [5, §3]. In [5, Theorem 1.1], Brochard verifies de Smit’s conjecture by proving the stronger statement that if $A,B$ are noetherian local rings with $\operatorname{edim}A\geq\operatorname{edim}B$ and $N$ is a finitely generated $B$-module that is flat over $A$, then $N$ is flat over $B$. In [6], Brochard, Khare, and I generalize, with a simpler proof, Brochard’s result as follows.
@@
Suppose $\varphi\colon A\to B$ is a local homomorphism of noetherian local rings, and $N$ is a nonzero finitely generated $B$-module whose flat dimension over $A$ satisfies $\operatorname{flat\,dim}_{A}N\leq\operatorname{edim}A-\operatorname{edim}B$, then $N$ is free as a $B$ module and the map $\varphi$ is complete intersection with $\operatorname{edim}A-\dim A=\operatorname{edim}B-\dim B$.
@@
The more general hypothesis in our theorem, as opposed to [5, Theorem 1.1], gives us the freedom to reduce its proof to the surjective case, using standard results in commutative algebra. This is instrumental in leading to a proof that is simpler than the one presented in [5] as we can apply an induction on $\operatorname{edim}A-\operatorname{edim}B$. A key ingredient in this induction step is the following theorem of Nagata’s.
@@
Question 2.5 has a positive answer when $A/{\mathfrak{m}}_{A}B$, the closed fiber of the map $A\to B$, is a complete intersection ring; this applies, in particular, when $A\to B$ is surjective.
In [7, Theorem 6.1] we are able to more prove more that what I stated above: When in addition the differential on $F$ satisfies $d(F)\subseteq{\mathfrak{m}}_{A}F$ (and this can always be arranged) there are also equalities
@@
where $b$ is the minimal number of generators of the $B$-module $\operatorname{H}_{0}(F)$. The binomial coefficients on the right are precisely the ranks of the free modules appearing in a Koszul complex on $n$-elements. This is reminiscent of phenomena well-known to commutative algebraists: Invariants attached to Koszul complexes often give good (typically, lower) bounds for general finite free complexes. Consider, for instance, the New Intersection Theorem [31], the ranks conjectures of Buchsbaum and Eisenbud [9], and the Total Rank Conjecture, now a theorem proved by Walker [35], and VandeBogert and Walker [34].
@@
Khare’s question also suggests a systematic study of invariants attached to complexes with derived actions of commutative rings. For instance: What restrictions arise on the, say, the Betti numbers, of a finite free $A$-complex $F$ if it admits a derived $B$-action? See [19, Lemma 10.7] for one such restriction.
@@
As noted above, in the context of modularity lifting, the goal is to prove that some $A$-module $M$ is free, or at least has a non-zero free summand, which suffices for many applications. To that end, one considers the conormal module of the map $\lambda$, namely the $\mathcal{O}$-module ${\mathfrak{p}}/{\mathfrak{p}}^{2}$, and, for any finitely generated $A$-module $M$, the congruence module of $M$, defined to be
@@
Here $M[I]$ denotes the submodule $\{m\in M\mid I\cdot m=0\}$ consisting of elements of $M$ annihilated by $I$. Number theoretic origins and motivations for this definition are given in the Introduction to [20]. The hypothesis that $\lambda$ is a regular point of codimension 0 ensures that both $\mathcal{O}$-modules ${\mathfrak{p}}/{\mathfrak{p}}^{2}$ and $\Psi_{\lambda}(M)$ have finite length; this is easy to check, by localizing at $\mathfrak{p}$}. It is equally easy to check that also the converse holds: If for some $\mathcal{O}$-valued point $\lambda$, either ${\mathfrak{p}}/{\mathfrak{p}}^{2}$ or $\Psi_{\lambda}(A)$ has finite length, then $\lambda$ is regular of codimension 0. See [14] or [20, Example 2.7] for examples arising from number theory.
@@
Number theoretic applications of the numerical criterion of Diamond, Lenstra, and Wiles rely on ingredients (like Ihara’s Lemma) that proved hard to generalize. More fundamental than this pragmatic difficulty is the point that the numerical criterion proves isomorphisms of complete intersections, and in later developments of Wiles’ methods the rings are not expected to be complete intersections. They are also situations where the rings are torsion and hence admit no $\mathcal{O}$-valued points. Applying a numerical criterion to patched rings circumvents these difficult as these are expected to be complete intersections, with sufficiently many regular $\mathcal{O}$-valued points, in more complex cases as well. However, the points are no longer of codimension zero, so one need extensions of the theory of congruence modules and Wiles defect that apply to points of higher codimension. This is achieved in [19]. What follows is an overview of our findings in op. cit., and also of some ongoing projects that build on it.
@@
The torsion submodule of conormal module ${\mathfrak{p}}/{\mathfrak{p}}^{2}$ (rather, its length) plays an important role in the sequel. Observe that the length of $\operatorname{tors}({\mathfrak{p}}/{\mathfrak{p}}^{2})$ equals the length of $\mathcal{O}/\operatorname{Fitt}_{c}({\mathfrak{p}}/{\mathfrak{p}}^{2})$. It follows from the Jacobi-Zariski sequence arising from the diagram $\mathcal{O}\to A\xrightarrow{\lambda}\mathcal{O}$ that there is a natural isomorphism
@@
where $\hat{\Omega}_{A/\mathcal{O}}$ is the universally finite module of differentials; see [25, §11]. Hence $\operatorname{Fitt}_{c}({\mathfrak{p}}/{\mathfrak{p}}^{2})$ equals $\operatorname{Fitt}^{A}_{c}(\hat{\Omega}_{A/\mathcal{O}})\cdot\mathcal{O}$, the $c$’th Fitting ideal of the module of Kähler differential extended to $\mathcal{O}$ along $\lambda$. The relevance of this observation is made clear in [20, §2.9].
@@
There is a closely related notion of a congruence ideal attached to $M$; please see [20] for details, including a description of the map $\eta_{\lambda}(M)$ in term of Yoneda extensions, and for the historical antecedents for these notion. In this paper, I will focus mainly on the following invariant
@@
  • The $M$ is free at any regular $\mathcal{O}$-valued point $\lambda$.
@@
Here ${\mathfrak{p}}^{(2)}$ is the second symbolic power of $\mathfrak{p}$}, namely ${\mathfrak{p}}^{2}A_{\mathfrak{p}}\cap A$. This theorem is proved by verifying that both components in the definition Wiles defect (3.4) change by the same amount, namely the order of the class of $f$, in the conormal module of $\lambda_{A}$; see [19, §8] and, for slightly different perspective, [21, §2].
@@
Fix $(A,\lambda)$ in $\operatorname{C}_{\mathcal{O}}(c)$, and suppose also that $A$ is Gorenstein. Let $M$ be an ${\mathfrak{m}}_{A}$-adically complete with $0pt_{A}M=\dim A$. If the map $\kappa_{\lambda}(M)$ from (3.5) is surjective, then $M\cong F\oplus W$ where $F$ is flat and $\mathfrak{p}$} is not in the support of $W$.
An module $M$ with the property that $0pt_{A}M=\dim A$ is said to be big Cohen-Macaulay module. It is well-understood that completion is often a good substitute for finite generation; Nakayama’s Lemma holds, for instance. However, we have not been able to verify the statement above by mimicking the argument for finitely generated modules given in [19, Theorem 9.2]. Instead, the proof we found uses in an essential way the theory of derived completions initiated by Greenlees and May [18], and is more satisfactory than the one in [19], for it gives an explicit splitting of $M$.
@@
The statement is agnostic to the values of the two invariants above. Thus, if one could construct a defect module the criteria for freeness (and complete intersection) described above can be phrased in terms of the vanishing of the defect module, and hence is sensible even when $M$ and the defect module are not finitely generated. Building on the discussion around Theorems 3.2 and 3.3, we have identified a potential candidate for the defect module: For $(A,\lambda)$ in $\operatorname{C}_{\mathcal{O}}(c)$ let $K$ be the Koszul complex on some finite generating set $a_{1},\dots,a_{n}$ for the ideal $\operatorname{Ker}\lambda$. With $c$ the codimension of $\lambda$, the defect module of a module $M$ (not necessarily finitely generated) should be the cokernel of the natural map
@@
induced by the action of $\operatorname{H}_{*}(K)$ on $\operatorname{H}_{*}(K\otimes_{A}M)$. This definition is also inspired by various criteria for detecting the complete intersection property of a local ring in terms of the multiplicative structure of the Koszul homology of the ring with respect to its maximal ideal, due to Assmus, Bruns, Tate, and Wiebe; see  Theorem 3.3 and also [8, §2.3]. We can verify, at least when both $A$ and $M$ has sufficient depth, that this defect module has the desired properties, and also that, when $M$ is in addition finitely generated, the length of the defect module is precisely $\delta_{\lambda}(M)$, the Wiles defect of $M$. Such a Koszul homology description of the defect module also makes it easier to compute the Wiles defect itself, in case $M$ is finitely generated. However, much remains to be done.

### v2

@@ Abstract
There are many excellent sources that give an overview of the work of Wiles’ proof; see, for example, [13, Chapter 1] and [14]. Roughly speaking, using Mazur’s [28] deformation theory of Galois representations, Wiles reduces the Shimura, Taniyama, Weil conjecture to proving that certain surjective maps $\varphi_{\Sigma}\colon R_{\Sigma}\to\mathbb{T}_{\Sigma}$ of (noetherian, commutative) local rings are isomorphisms. The ring $R_{\Sigma}$ parametrizes Galois deformations subject to constraints determined by a finite collection of primes $\Sigma$ and $\mathbb{T}_{\Sigma}$ is a Hecke algebra subject to the same constraints, which acts on a certain space of modular forms, $M_{\Sigma}$. To prove that $\varphi_{\Sigma}$ is an isomorphism, Wiles first proves that $M_{\Sigma}$ is free as a $\mathbb{T}_{\Sigma}$-module—this is a delicate argument—and also that $\mathbb{T}_{\Sigma}$ is a complete intersection. He then uses these properties to prove that $\varphi_{\Sigma}$ is an isomorphism; as a corollary, he deduces the freeness of $M_{\Sigma}$ also as an $R_{\Sigma}$-module. It is Diamond’s insight [15] that one can prove the latter statement directly and that this implies $\varphi_{\Sigma}$ is an isomorphism of complete intersections, and also that $M_{\Sigma}$ is free over $\mathbb{T}_{\Sigma}$.
@@
where the map on the left is a homomorphism of rings, say with kernel $I$, and $M_{\varnothing}\cong M_{\infty}/IM_{\infty}$. The point of the patching process is that $R_{\infty}$ is a regular local ring and $M_{\infty}$ is a maximal Cohen-Macaulay module over it. Thus the Auslander-Buchsbaum formula implies that $M_{\infty}$ is a free $R_{\infty}$-module, and hence that $M_{\varnothing}$ is a free $R_{\varnothing}$-module, as desired; see Proposition 2.1. It turns out that $I$ is generated by a regular sequence, so as a byproduct of the patching argument one gets that $R_{\varnothing}$ is a complete intersection; this is special to Wiles’ context and important for the passage from the minimal to the general, non-minimal, case. This passage is facilitated by a new criterion, discovered by Diamond [15] (building on work of Wiles and Lenstra [26]) that detects when a finitely generated module over a local ring is free and the ring is a complete intersection; see Theorem 3.1.
@@
Commutative algebra thus enters in both steps of the proof. As such, the result that Wiles and Diamond wanted (the Auslander-Buchsbaum formula) had been available since the late 1960’s. In contrast, to tackle the non-minimal case, Wiles and Diamond needed to develop new commutative algebra, and in particular a numerical criterion (in terms that are meaningful in number theory) for detecting freeness of modules over local rings. Nevertheless, each step raises new questions, and leads to new results, in commutative algebra. This is what is described below.
@@
The patching method is (or so it seems to me) specific to the number theoretic context. In Section 2, I discuss the commutative algebra that is partly motivated by a desire to simplify, and even entirely do away with, the patching argument of Taylor and Wiles to establish modularity in the minimal case. It starts with a conjecture of de Smit (now a theorem of Brochard’s) that, in hindsight, can be interpreted as a converse to a well-known (to commutative algebraists!) result of Nagata on a special kind of surjective complete intersection map [^1] However, the hypothesis of the de Smit conjecture provides the seed of a patching datum that develops into full fledged patching. The patching argument has become an indispensable tool in the number theorist’s toolkit: it introduces a useful rigidity and robustness, allowing one to, for example, invert the prime $p$ when proving modularity of $p$-adic Galois representations, which the hypothesis of de Smit’s conjecture does not provide..
@@
The patching method of Taylor and Wiles has been, and continues to be, developed extensively, and is now available also in the non-minimal case, thanks to work of Kisin [24], and this has often allowed one to bypass the numerical criterion. However, although this has led to much more general modularity lifting theorems, among other applications, it has been at the expense of some refined torsion information; all this is explained in the introduction to [19]. See also the discussion after Theorem 3.3. The main contribution of op. cit. is to establish a numerical criterion that applies in greater generality, so that it can be applied to rings and modules after patching.
@@
In what follows, I assume familiarity with basic concepts from commutative algebra; much of what is needed can be found in Chapter 1 of the book by Bruns and Herzog [8], and I take that as the standard reference. The homological algebra notions that come up are already present, for most parts, in the book by Cartan and Eilenberg [12]. Given a local ring $A$, I write ${\mathfrak{m}}_{A}$ for its maximal ideal and $k_{A}$ for its residue field $A/{\mathfrak{m}}_{A}$. The embedding dimension of A, denoted $\operatorname{edim}A$, is the $k_{A}$-rank of the Zariski tangent space ${\mathfrak{m}}_{A}/{\mathfrak{m}}_{A}^{2}$; equivalently, the minimal number of generators of the ideal ${\mathfrak{m}}_{A}$.
@@
Subsequent to Diamond’s work, Bart de Smit conjectured that if $A\to B$ is a local homomorphism of artinian local rings with $\operatorname{edim}A=\operatorname{edim}B$, then any $B$-module that is flat as an $A$-module is also flat as a $B$-module. This strengthens Proposition 2.1 in the case where $A$ and $B$ have the same embedding dimension, and allows one to dispense with the patching technique in the work of Wiles; see [5, §3]. In [5, Theorem 1.1], Brochard verifies de Smit’s conjecture by proving the stronger statement that if $A,B$ are noetherian local rings with $\operatorname{edim}A\geq\operatorname{edim}B$ and $N$ is a finitely generated $B$-module that is flat over $A$, then $N$ is flat over $B$. In [6], Brochard, Khare, and I generalize, with a simpler proof, Brochard’s result as follows.
@@
Suppose $\varphi\colon A\to B$ is a local homomorphism of noetherian local rings, and $N$ is a nonzero finitely generated $B$-module whose flat dimension over $A$ satisfies $\operatorname{flat\,dim}_{A}N\leq\operatorname{edim}A-\operatorname{edim}B$, then $N$ is free as a $B$-module and the map $\varphi$ is complete intersection with $\operatorname{edim}A-\dim A=\operatorname{edim}B-\dim B$.
@@
The more general hypothesis in our theorem, as opposed to [5, Theorem 1.1], gives us the freedom to reduce its proof to the surjective case, using standard results in commutative algebra. This is instrumental in leading to a proof that is simpler than the one presented in [5] as we can apply an induction on $\operatorname{edim}A-\operatorname{edim}B$. A key ingredient in this induction step is the following theorem of Nagata.
@@
Question 2.5 has a positive answer when $B/{\mathfrak{m}}_{A}B$, the closed fiber of the map $A\to B$, is a complete intersection ring; this applies, in particular, when $A\to B$ is surjective.
In [7, Theorem 6.1] we prove more than what I stated above: When in addition the differential on $F$ satisfies $d(F)\subseteq{\mathfrak{m}}_{A}F$ (and this can always be arranged) there are also equalities
@@
where $b$ is the minimal number of generators of the $B$-module $\operatorname{H}_{0}(F)$. The binomial coefficients on the right are precisely the ranks of the free modules appearing in a Koszul complex on $n$-elements. This is reminiscent of phenomena well-known to commutative algebraists: Invariants attached to Koszul complexes often give good (typically, lower) bounds for general finite free complexes. Consider, for instance, the New Intersection Theorem [31], the rank conjectures of Buchsbaum and Eisenbud [9], and the Total Rank Conjecture, now a theorem proved by Walker [35], and VandeBogert and Walker [34].
@@
Khare’s question also suggests a systematic study of invariants attached to complexes with derived actions of commutative rings. For instance: What restrictions arise on, say, the Betti numbers of a finite free $A$-complex $F$ if it admits a derived $B$-action? See [19, Lemma 10.7] for one such restriction.
@@
As noted above, in the context of modularity lifting, the goal is to prove that some $A$-module $M$ is free, or at least has a nonzero free summand, which suffices for many applications. To that end, one considers the conormal module of the map $\lambda$, namely the $\mathcal{O}$-module ${\mathfrak{p}}/{\mathfrak{p}}^{2}$, and, for any finitely generated $A$-module $M$, the congruence module of $M$, defined to be
@@
Here $M[I]$ denotes the submodule $\{m\in M\mid I\cdot m=0\}$ consisting of elements of $M$ annihilated by $I$. Number theoretic origins and motivations for this definition are given in the Introduction to [20]. The hypothesis that $\lambda$ is a regular point of codimension 0 ensures that both $\mathcal{O}$-modules ${\mathfrak{p}}/{\mathfrak{p}}^{2}$ and $\Psi_{\lambda}(M)$ have finite length; this is easy to check, by localizing at $\mathfrak{p}$}. It is equally easy to check that also the converse holds: If for some $\mathcal{O}$-valued point $\lambda$, either ${\mathfrak{p}}/{\mathfrak{p}}^{2}$ or $\Psi_{\lambda}(A)$ has finite length, then $\lambda$ is regular of codimension 0. See [14, §5.1] or [20, Example 2.7] for examples arising from number theory.
@@
Number theoretic applications of the numerical criterion of Diamond, Lenstra, and Wiles rely on ingredients (like Ihara’s Lemma) that proved hard to generalize. More fundamental than this pragmatic difficulty is the point that the numerical criterion proves isomorphisms of complete intersections, and in later developments of Wiles’ methods the rings are not expected to be complete intersections. There are also situations where the rings are torsion and hence admit no $\mathcal{O}$-valued points. Applying a numerical criterion to patched rings circumvents these problems as these are expected to be complete intersections, with sufficiently many regular $\mathcal{O}$-valued points, in more complex cases as well. However, the points are no longer of codimension zero, so one need extensions of the theory of congruence modules and Wiles defect that apply to points of higher codimension. This is achieved in [19]. What follows is an overview of our findings in op. cit., and also of some ongoing projects that build on it.
@@
The torsion submodule of the conormal module ${\mathfrak{p}}/{\mathfrak{p}}^{2}$ (rather, its length) plays an important role in the sequel. Observe that the length of $\operatorname{tors}({\mathfrak{p}}/{\mathfrak{p}}^{2})$ equals the length of $\mathcal{O}/\operatorname{Fitt}_{c}({\mathfrak{p}}/{\mathfrak{p}}^{2})$. It follows from the Jacobi-Zariski sequence arising from the diagram $\mathcal{O}\to A\xrightarrow{\lambda}\mathcal{O}$ that there is a natural isomorphism
@@
where $\hat{\Omega}_{A/\mathcal{O}}$ is the universally finite module of differentials; see [25, §11]. Hence $\operatorname{Fitt}_{c}({\mathfrak{p}}/{\mathfrak{p}}^{2})$ equals $\operatorname{Fitt}^{A}_{c}(\hat{\Omega}_{A/\mathcal{O}})\cdot\mathcal{O}$, the $c$’th Fitting ideal of the module of Kähler differentials extended to $\mathcal{O}$ along $\lambda$. The relevance of this observation is made clear in [20, §2.9].
@@
There is a closely related notion of a congruence ideal attached to $M$; please see [20] for details, including a description of the map $\eta_{\lambda}(M)$ in term of Yoneda extensions, and for the historical antecedents for these notions. In this paper, I will focus mainly on the following invariant
@@
  • The module $M$ is free at any regular $\mathcal{O}$-valued point $\lambda$.
@@
Here ${\mathfrak{p}}^{(2)}$ is the second symbolic power of $\mathfrak{p}$}, namely ${\mathfrak{p}}^{2}A_{\mathfrak{p}}\cap A$. This theorem is proved by verifying that both components in the definition of the Wiles defect (3.4) change by the same amount, namely the order of the class of $f$, in the conormal module of $\lambda_{A}$; see [19, §8] and, for slightly different perspective, [21, §2].
@@
Fix $(A,\lambda)$ in $\operatorname{C}_{\mathcal{O}}(c)$, and suppose also that $A$ is Gorenstein. Let $M$ be an ${\mathfrak{m}}_{A}$-adically complete $A$-module with $0pt_{A}M=\dim A$. If the map $\kappa_{\lambda}(M)$ from (3.5) is surjective, then $M\cong F\oplus W$ where $F$ is flat and $\mathfrak{p}$} is not in the support of $W$.
An $A$-module $M$ with the property that $0pt_{A}M=\dim A$ is said to be a big Cohen-Macaulay module. It is well-understood that completion is often a good substitute for finite generation; Nakayama’s Lemma holds, for instance. However, we have not been able to verify the statement above by mimicking the argument for finitely generated modules given in [19, Theorem 9.2]. Instead, the proof we found uses in an essential way the theory of derived completions initiated by Greenlees and May [18], and is more satisfactory than the one in [19], for it gives an explicit splitting of $M$.
@@
The statement is agnostic to the values of the two invariants above. Thus, if one can construct a defect module the criteria for freeness (and complete intersection) described above can be phrased in terms of the vanishing of the defect module, and hence is sensible even when $M$ and the defect module are not finitely generated. Building on the discussion around Theorems 3.2 and 3.3, we have identified a potential candidate for the defect module: For $(A,\lambda)$ in $\operatorname{C}_{\mathcal{O}}(c)$ let $K$ be the Koszul complex on some finite generating set $a_{1},\dots,a_{n}$ for the ideal $\operatorname{Ker}\lambda$. With $c$ the codimension of $\lambda$, the defect module of a module $M$ (not necessarily finitely generated) should be the cokernel of the natural map
@@
induced by the action of $\operatorname{H}_{*}(K)$ on $\operatorname{H}_{*}(K\otimes_{A}M)$. This definition is also inspired by various criteria for detecting the complete intersection property of a local ring in terms of the multiplicative structure of the Koszul homology of the ring with respect to its maximal ideal, due to Assmus, Bruns, Tate, and Wiebe; see  Theorem 3.3 and also [8, §2.3]. We can verify, at least when both $A$ and $M$ have sufficient depth, that this defect module has the desired properties, and also that, when $M$ is in addition finitely generated, the length of the defect module is precisely $\delta_{\lambda}(M)$, the Wiles defect of $M$. Such a Koszul homology description of the defect module also makes it easier to compute the Wiles defect itself, in case $M$ is finitely generated. However, much remains to be done.